Closed Actions Booster Commissioning Working Group

Assigned to Start date Description State Result
M.Eshraqi 2007-12-04 Check feasibility to move quadrupoles in LTB line to reduce gradients. It was decided to leave the quadrupoles at their current position and rather buy new power supplies in case this is needed. edit
K.Hanke 2007-12-04 Prepare complete list of all magnets in transfer and measurement lines between LINAC4 and PSB. Done. This action has been taken over by A. Lombardi. edit
T.Zickler 2007-12-04 Verify suitability of LT.BHZ20 parameters with LINAC4 operation. Increase water pressure to 10 bars (magnet) and replace power converter. edit
C.DeAlmeida 2007-12-04 Cross-check parameters of all power supplies (see list above) if suitable for LINAC4 operation. Done. edit
K.Hanke 2007-12-04 Identify responsibles for the kickers in the LBE line. Action has been transferred to the BT group. edit
T.Zickler 2007-12-04 Verify the quadrupole parameters in the LBE line in view of LINAC4 operation. The quadrupole parameters could be fixed after the accepted upgrade proposals for LBE and LBS line (see 2010 minutes). edit
T.Zickler, C.DeAlmeida 2008-01-07 Provide A. Lombardi with information concerning the operational range of the complete equipment of the old line. Done. See minutes of March 20th. edit
A.Lombardi 2008-01-07 Update magnet inventory for LINAC4 and LINAC4-to-PSB transfer line including the measurement lines and make document public on EDMS. OK. edit
W.Weterings 2008-02-07 Clarify with C. De Almeida if it would be possible to operate BI.SMV and BI.BVT with the new power supplies and Linac2 operation. See also minutes from 20/03/2008. The power supplies should work for both Linac4 and Linac2 operation. edit
K.Hanke 2008-02-07 Contact RP to find out if it is acceptable to have the beam dumped continuously on a beam stopper during the commissioning period. This is not possible. edit
K.Hanke 2008-02-07 Identify with R. Scrivens a possible location for a beam stopper at the end of the Linac2 line for the transfer line commissioning period. Irrelevant as it is not acceptable to use a beam stopper for commissioning. edit
C.DeAlmeida 2008-02-07 Clarify with S. Baird if the consolidation budget can cover the replacement of the old Linac2 Tekelec power supplies. See minutes 10/04/2008. edit
A.Lombardi, M.Eshraqi 2008-03-20 Check aperture limitations of BI line both for Linac4 and Linac2 operation. In case the aperture would be too stringent, inform W.Weterings as soon as possible. See presentation 10/04/2008. edit
W.Weterings 2008-03-20 Can the installation of some new PSB injection elements be advanced? Closed. See minutes 10/04/2008. edit
K.Hanke 2008-03-20 Draft a modification to the planning of Linac4 transfer line commissioning and PSB injection line changes. Done. See minutes of July 3rd 2008. edit
T.Fowler, W.Weterings 2008-03-20 Clarify who within BT could take over the responsibility for the kicker magnets of the LBE line. W. Weterings agreed to act as contact person for this equipment. edit
K.Hanke 2008-03-20 Perform simulations of the LBE line with Linac4 to check suitability of magnets and power supplies. The LBE line has been redesigned with a different approach for the emittance measurement (sLHC-PROJECT-Report-0029). edit
W.Weterings 2008-04-10 Contact T. Fowler to find out if the installation of BI.KSW could be advanced. It seems feasible to advance the BI.KSW installation to 2011/2012 shutdown. edit
J.Borburgh 2008-04-10 Review design of the new septum. Contact T. Fowler to find out whether the kick in the distributor could be increased. See minutes 22/05/2008. edit
M.Eshraqi 2008-04-10 Provide simulations to decide if the horizontal aperture of the septum could further be reduced from 70 to 50 mm. Done. This is in principle possible, but detailed error studies still have to be done. edit
C.DeAlmeida 2008-04-10 Correct EVM document to include all agreed modifications of the power converters in the Linac4 to PSB transfer line and inform M. Vretenar about the changes. Done. edit
K.Hanke 2008-04-10 Find out if collimators could be added upstream of the septum. This action has been attributed to the Linac4 beam dynamics working group. edit
K.Hanke 2008-05-26 Obtain specifications of kicker magnets in LBE line and send them to W. Weterings. Done. edit
T.Zickler, C.DeAlmeida 2008-05-26 Check if the current BI.DVT equipment is adapted to allow for an increase in deflection angle by a factor 2.4. Has become irrelevant as injection chicane layout has again been changed. edit
K.Hanke 2008-05-26 Contact S. de Man to organise the removal of the lead distributor for the 2011/2012 shutdown. Will be done during 2009/2010 shutdown (still to be organised). edit
K.Hanke, W.Weterings 2008-11-18 Access PSB to check if PSB injection dump could be located in the prolongation of the line behind the wall. PSB injection region has been visually inspected by W. Weterings and B. Mikulec on 19th of March 2009 (photographs taken by W. Weterings). The layout of the region has to be found. edit
T.Hermanns 2009-01-20 Provide specifications for beam load on LBE/LBS line dumps once a decision has been made on the line upgrades. EDMS Document 1056611. edit
A.Lombardi 2009-01-20 Identify appropriate positions for Linac2 line collimators to minimize activation of distributor due to halo particles. It is impossible due to space restrictions to place collimators in the Linac4 <160 MeV part. It might be possible in the Linac4/Linac2 transfer line, but first the damage potential to the PSB injection equipment should be proven. edit
K.Hanke, B.Mikulec 2009-06-22 Discuss with PO how to implement Linac4 power converter interlocks in the BIS framework. See EDMS document 1016233. edit
D.Kuechler, R.Scrivens 2009-06-22 Identify the most efficient way to cut the Linac4 beam rapidly at the source level (3 MeV test stand). The BIS should send a signal to set the screen grid voltage on the source to 0 V (~50 us reaction time) and in parallel a logic pulse to the LL RF system (see minutes 7/12/2009). edit
K.Hanke, B.Mikulec 2009-06-22 Obtain final input from the Linac4 colleagues to provide BLM specifications to BI. Done. The specifications have been sent to B. Dehning. edit
L.Soby 2009-06-22 Can the fast transformers in 8L1 be connected to the watchdog system? See Minutes 21/09/2009. edit
B.Mikulec 2009-06-22 Write an EDMS document defining the interlock specifications. EDMS document 1016233 has been published. edit
U.Raich 2009-07-30 Discuss with E. Bravin the aspect of electron stripping in scintillating screens and find somebody to study this effect. BI have made simulations and will soon publish an internal note. edit
T.Hermanns 2009-09-21 Check if 30 mA of beam current are sufficient to obtain and energy measurement with the required precision in the LBS line. Yes, 30 mA are sufficient (see sLHC note on LBS line). edit
J.Tan 2009-09-21 Follow up with L. Soby and P. Odier if it would be possible to use the 'middle-fast' transformer installed in PSB section 8L1 for injection efficiency measurements of the Linac4 beam. It is possible to use the intermediate frequency BCTs for this purpose. The electronics has to be upgraded. edit
B.Dehning 2009-09-21 Clarify sensitivity of Linac4 BLMs and the flexibility of sensitivity selection. See minutes of the Linac4 diagnostics working group from 14/12/2009. edit
A.Lombardi 2009-09-21 Discuss with F.Gerigk the potential of the debuncher to increase the energy spread and assess the usefulness of 2 cavities. It has been decided to install only 1 cavity at a new position. The original position of the cavity will be kept as backup in case the amplifier would not be compatible with the parameters at the new position. edit
F.Lenardon, L.Soby 2009-10-05 Estimate the time required to send the information on the Linac4 beam current to the interlock system. The beam pulse will be sampled at 10 MHz. With the current system it would be possible to interlock the source for the following beam pulse. However this interlock would be done purely by software and is therefore not safe (this is similar to the Linac-2 watchdog). Should a hardware interlock be needed then this must be specified and discussed. As a different solution one could install TRIC cards with digital watch dogs and transform the signal to optical. To be seen. edit
K.Hanke 2009-10-07 Evaluate the possibility to install an internal dump in the PSB. Not further pursued. edit
B.Mikulec 2009-12-07 Follow up a separate interlock can be created for the BSW and connected to the BIC. This will be possible. edit
A.Newborough 2010-02-07 Follow-up the upgrade of the SMIT-Air quadrupoles in the LBE line (fabrication of new coils, installation of water cooling) and make sure that this project is added to the magnet work package. Contact S. Baird to find out if a part of the cost for the coil fabrication can be covered by the consolidation budget. See minutes 11/03/2010. Information forwarded to T. Zickler. Discussion with S. Baird delayed until decision on PSB 2 GeV upgrade. edit
T.Hermanns 2010-02-07 Obtain information concerning the transport facilities for LBE/LBS. See minutes 11/03/2010. edit
B.Mikulec 2010-02-07 Contact S. Deleval to start discussions about the connection of the water circuits. See minutes 11/03/2010. This should be followed up by the BCC. edit
K.Hanke 2010-02-25 Check if power converters between LTB.BHZ30 and LTB.BHZ40 can be bi-polarly operated in ppm mode. A. Lombardi replied that she believes that ions can be transported to the LBE/LBS lines just by recomputing the optics with the triplet in the ITH line. edit
T.Hermanns 2010-02-25 The power converters for the LBE line magnets should be bi-polar. Inform and discuss again with D. Nisbet and J.M. Cravero. The discussion took place and an email with all requested specifications for the power supplies has been sent to D. Nisbet and J-M. Cravero on 12/03/2010. edit
R.Chamizo 2010-03-16 Initiate discussions with RP and BI concerning the LBS line dump. Re-iterate simulations if modifications are needed for the dump design. This will be followed up by Cesare Maglioni. edit

-- BettinaMikulec - 04 Dec 2007

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2008-06-30 - FrankGerigk
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    SPL All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright